OpenStudy (anonymous):

You are rolling two 6-sided dice and examining the sum of the two dice (for example, if you roll a 3 and a 5, this is an 8). a) What is the probability of rolling a sum of 7? b) What is the probability of rolling a sum of 11? c) What is the probability of rolling a sum of 7 or 11? d) What is the probability of rolling a sum of 7 and 11?

OpenStudy (anonymous):

probability of rolling 7 is 1/6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the way i do these is make a table that is 6x6

OpenStudy (anonymous):

then add the corresponding terms...so it would look like this 1 2 3 4 5 6 7 8 1 2 3.... 2 3 3 4............. 4 5 6 7 8

OpenStudy (anonymous):

Ok i see!:) what about the rest

OpenStudy (anonymous):

add them all up

OpenStudy (anonymous):

and for the probability of summing a 7...would be counting the number of 7's when you add it all up. divided by 36

OpenStudy (anonymous):

so a i got 1/6 right?

OpenStudy (anonymous):

where do i start for b?

OpenStudy (anonymous):

1 2 3 4 5 6 1 2 3 4 5 6 7 2 3 4 5 6 7 8 3 4 5 6 7 8 9 4 5 6 7 8 9 10 5 6 7 8 9 10 11 6 7 8 9 1011 12

OpenStudy (anonymous):

so a is 7/36?

OpenStudy (anonymous):

thats what it should look like...excluding the first row and first column. count up the number of 7's along the diagnol. And get 6 so 6/36=1/6

OpenStudy (anonymous):

thats what i meant sorry ok i get a

OpenStudy (anonymous):

yes and 11 would be 2/36

OpenStudy (anonymous):

ok ya

OpenStudy (anonymous):

ignore what i typed earlier..it does not go up to 8. Its a 6 sided die

OpenStudy (anonymous):

are you in AP statistics?

OpenStudy (anonymous):

now what do i do for c and d

OpenStudy (anonymous):

yes. and so am i wrong

OpenStudy (anonymous):

(a) 1/6 (b) 2/36 (c)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

for or you add. and for and you multiply

OpenStudy (anonymous):

and for c?

OpenStudy (anonymous):

its 3/36 or 1/12?

OpenStudy (anonymous):

because the trials are indpendent, add 6/36 + 2/36 and get 8/36

OpenStudy (anonymous):

ok and d?

OpenStudy (anonymous):

for d) the trials are independent so you can multiply the probabilities: 1/6 (2/36)=2/216

OpenStudy (anonymous):

ok thanjs!

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

and A card is drawn from a standard deck of 52 cards (13 cards in each of four different suits). What is the probability that the card is a heart (one of the four suits)? would be 0.25 right??

OpenStudy (anonymous):

it would help alot if you made me a fan :P

OpenStudy (anonymous):

yes